### v1

# Diagnosing phase transitions through time scale entanglement

Abstract
We propose a new method based on the compression of arbitrary multi-point correlators into quantics tensor trains and studying its so called bond dimension as an indicator of time scale entanglement in the system. This quantics tensor train diagnostics approach allows us to i) identify quantum phase transitions (ground state crossings) and thermal crossovers, ii) distinguish between these two by allowing flexibility through iii) using arbitrary available correlators and iv) the possibility to deal with low-accuracy input data through varying the quantics tensor train cutoffs. We verify the validity of the approach by studying the phase diagram of the Hubbard dimer, the four-site Hubbard ring with and without next-nearest neighbor hopping and the single-impurity Anderson model. We find time scale entanglement to be a system-inherent property becoming maximal at phase transitions and crossover, which translates to peaks in the quantics tensor train bond dimension of correlator functions. Additionally, the connection of the method to other physical quantities, such as various entanglement measures and susceptibilities, is shown.
preprint: APS/123-QED
I Introduction
Phase transitions are among the most intriguing phenomena in physics. Yet, their numerical treatment, particularly in systems with nontrivial electronic behavior, can be challenging, especially when exotic phases emerge [1, 2, 3]. In particular, the full wave function is not available for many systems, and using susceptibilities to diagnose transitions can be brittle, since we need to know a priori where to look, i.e., we must make sure to compute the response associated with the transition in question or at least one that is not insensitive to the transition. In general, we can expect some but a much weaker signal in the susceptibilities that are not the one associated with the symmetry breaking for a second order phase transition. A first order phase transitions, on the other hand, will not be signaled at all in the susceptibilities prior to the phase transition.
Tensor trains, used originally as matrix product states (MPSs) for the ground state of spin systems [4, 5], have been elevated in the last years to a general compression ansatz [6, 7, 8]. By combining tensor trains with the quantics representation – a set of auxiliary qubits for exponentially different time or length scales – one obtains a compression ansatz for arbitrary functions in space and time [9, 10, 11] called a quantics tensor train (QTT). QTTs were instrumental, among others, in computational breakthroughs in modeling of turbulence [12, 13, 14, 15, 16], plasma physics [17], and quantum chemistry [18].
Some of the authors of this Letter have recently extended the use of QTTs for modeling electronic correlation functions [19], and found that it unifies prior pole [20, 21], basis [22, 23, 24, 25], and mesh refinement [26] ansätze into a single framework – one where doubling the size of the system amounts to adding a qubit, a mere linear increase. When combined with a way to sample QTTs [27, 28], this framework allows the solution of diagrammatic theories for systems which were previously inaccessible [29]. Moreover, full $n$-particle correlators can be compressed and stored.
The compression power of tensor trains depends on the connection strength between the qubits and thus between the length or time scales, encoded in the so-called bond dimension $D_{\mathrm{max}}$. For physical qubits used in MPSs for spin chains, this bond dimension has physical significance: it is related to the entanglement between two parts of the system [30]; for auxiliary qubits used in QTTs for correlation functions, it has until now only been used as a performance metric.
Figure 1: QTT bond dimension $D_{\mathrm{max}}$ of different imaginary-time correlation functions for the Hubbard dimer as a function of electronic repulsion $U$ and chemical potential $\mu/U$: (a) $D_{\mathrm{max}}$ for the one-particle Green’s function $G_{11}^{\uparrow\uparrow}$ at temperature $\beta^{-1}=1/50$, and (b,d) for the two-particle Green’s function $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}$ at $\beta^{-1}=1/50$ and $1/30$, respectively. For comparison, the value of the static, local spin susceptibility $\chi_{s}$ at $\beta^{-1}=1/50$ is shown in panel (c). Red dashed lines mark crossings of the ground state of the model. These crossings can be diagnosed as a sharp maximum in $D_{\mathrm{max}}$.
In this Letter, we argue that the bond dimension of correlation functions has physical significance: it becomes maximal at crossovers and phase transitions [red dots in Fig. 1]. These maxima appear in both single- [Fig. 1 (a)] and two-particle propagators [Fig. 1 (b,c)], without the need to identify the specific susceptibility associated with the transition. We call this program of detecting phase transitions quantics tensor train diagnostics (QTTD).
II Quantics tensor trains
In the QTT approach [6, 19] each variable is represented through a set of binary numbers or “quantics” corresponding to different length or time scales. The resulting tensor is then factorized into a tensor train (TT) at each scale. To illustrate the approach, let us consider a function $f(\tau)$ of a discretized variable $\tau$, with $\tau\in\{0,\ldots,M-1\}$ on $M=2^{R}$ grid points. In the quantics approach, $\tau$ is expressed in its binary representation $\tau=(\sigma_{1}\sigma_{2}\dots\sigma_{R})_{2}=\sum_{\ell=1}^{R}2^{R-\ell}\sigma_{\ell},$ with $\sigma_{\ell}\in\{0,1\}\,$ leading to $f(\tau)$ being seen as a $2\times 2\times...\times 2$ ($R$ times) tensor $F_{\sigma_{1},\ldots,\sigma_{R}}$ instead. Now each tensor index $\sigma_{\ell}$ corresponds to an exponentially distinct length or time scale of the system. The first bit $\sigma_{1}$ represents the coarsest scale which divides the system into halves, while the last bit $\sigma_{R}$ reflects the finest scale.
Factorization of this tensor into a (truncated) tensor train (TT) or matrix product state (MPS) of the form
$F_{\sigma_{1}\ldots\sigma_{R}}\approx\tilde{F}_{\sigma_{1}\ldots\sigma_{R}}=\!\sum_{\alpha_{1}=1}^{D_{1}}\!\ldots\!\sum_{\alpha_{R-1}=1}^{D_{R-1}}[M_{1}]^{\sigma_{1}}_{1\alpha_{1}}[M_{2}]^{\sigma_{2}}_{\alpha_{1}\alpha_{2}}\cdots[M_{R}]^{\sigma_{R}}_{\alpha_{R-1}1}$ (1)
can be achieved by means of singular value decomposition (SVD) [19] or using the tensor cross interpolation (TCI) algorithm [8, 28, 27]. In Eq. (1), each $M_{\ell}$ is a $D_{\ell-1}\times 2\times D_{\ell}$ tensor with “physical” binary index $\sigma_{\ell}$ and virtual indices (“bonds”) $\alpha_{\ell-1},\alpha_{\ell}$, which are summed over. Eq. (1) translates to:
[Uncaptioned image] (2)
in tensor network form [5]. We define the (maximum) bond dimension of the QTT as $D_{\mathrm{max}}=\textrm{max}_{\ell}(D_{\ell})$. Generally, the bond dimensions $D_{\ell}$ characterize the amount of “length or time scale entanglement” in the system are truncated in the TT factorization such that a specified error cutoff or tolerance $\epsilon$ is satisfied. In the case of SVD, usually the squared Frobenius norm $\epsilon=||F-\tilde{F}||^{2}_{\mathrm{F}}/||F||^{2}_{\mathrm{F}}$ is taken as a measure of the error of the approximation in Eq. (1). Many problems in many-body physics have been shown to use quantities that are strongly compressible with the QTT approach, which is possible if the bond dimension is sufficiently small [19, 31, 32, 33, 28, 29, 34, 35, 36, 37].
III Quantics tensor trains diagnostics
Let us start from a general imaginary time two-point correlator $G_{AB}(\tau)=-\langle\hat{\mathcal{T}}\hat{A}(\tau)\hat{B}\rangle$ with the fermionic operators $\hat{A},\hat{B}$ and the time ordering operator $\hat{\mathcal{T}}$. Then, $G_{AB}(\tau)$ can be expressed in its Lehmann representation:
$G_{AB}(\tau)=-\frac{1}{Z}\sum_{n}e^{-\beta\epsilon_{n}}\sum_{m}e^{-\tau(\epsilon_{m}-\epsilon_{n})}\langle n|\hat{A}|m\rangle\langle m|\hat{B}|n\rangle,$ (3)
where $1/\beta$ is temperature, $\tau\in[0,\beta)$ is imaginary time, $Z$ is the partition function, $\hat{H}$ is the Hamiltonian including the chemical potential, and $\hat{H}|n\rangle=\epsilon_{n}|n\rangle$ defines its eigenenergies and eigenstates. As can be seen, $G_{AB}(\tau)$ is the sum of exponentials. A single exponential has a QTT representation of bond dimension one, since it is a product of exponentials of the binary variables $\sigma_{l}$ at each scale: $e^{\tau\epsilon}=\prod_{\ell=1}^{R}e^{\epsilon\beta 2^{-\ell}\sigma_{\ell}}$. The bond dimension of the exact QTT representation of $G_{AB}$ is, thus, bound by the number of nonzero elements in the sum. Let us now consider the case of temperature approaching zero ($\beta\rightarrow\infty$), then
$G_{AB}(\tau)$ $=-\frac{1}{N}\sum_{i=1}^{N}\sum_{m}\Big{(}e^{-\tau\epsilon_{m}}\langle\mathrm{GS}_{i}|\hat{A}|m\rangle\langle m|\hat{B}|\mathrm{GS}_{i}\rangle$
$\quad+e^{-(\beta-\tau)\epsilon_{m}}\langle m|\hat{A}|\mathrm{GS}_{i}\rangle\langle\mathrm{GS}_{i}|\hat{B}|m\rangle\Big{)},$ (4)
where $|\mathrm{GS}_{i}\rangle$ denotes the $i-$th ground state and $N$ is the number of ground states (states with the lowest energy). We set $\epsilon_{\mathrm{GS}}=0$; and the first (second) term contributes predominately for small (large) values of $\tau$.
First, let us focus on the case of a single non-degenerate ground state. Then, the bond dimension of the exact QTT representation is bound by the number of nonzero terms in the sum, so at maximum by $M-1$ (if all matrix elements are nonzero) for $M$ eigenstates $m$. Next, consider a ground state crossing of the two lowest states $|\mathrm{GS}_{1}\rangle$ and $|\mathrm{GS}_{2}\rangle$ ($N=2$). Then the bond dimension of the exact QTT is at most $22M-$ 2 since twice as many exponentials can contribute. In practice, it may be lower as only some of the states $|m\rangle$ will give nonzero matrix elements. Still, at the ground state crossing the sum consists of the exponentials and matrix elements associated with $|\mathrm{GS}_{1}\rangle$ and $|\mathrm{GS}_{2}\rangle$ leading to larger bond dimensions of $G_{AB}(\tau)$ than before or after the crossing. Therefore, at zero temperature, a quantum phase transition (QPT) or a ground state crossing of the Hamiltonian can be determined by monitoring the bond dimension of the QTT representation. We will call this approach quantics tensor train diagnostics (QTTD), where the diagnostics power of the approach is further examined in Appendix D. Similarly, ground state crossings of degenerate states (e.g. a doublet with a triplet state) will lead to maxima of the bond dimension in parameter space.
Let us now consider finite, but low temperatures. Then, also excited states $|n\rangle$ contribute to $G_{AB}(\tau)$, suppressed by $e^{-\beta\epsilon_{n}}$. However, if we compress the two-point correlator into a QTT, most of these contributions will be truncated, if they are below a certain specified cutoff in the QTT construction. Hence, we will still see a peak of the maximum bond dimension in the vicinity of the corresponding QPT (at $T=0$), which will become less sharp with increasing temperature which smears out this transition.
The QTTD approach can also be used to identify thermal crossovers of different regimes, such as the Kondo effect, that cannot be identified by merely looking at the lowest energy states in the ($T$-independent) eigenspectrum of the Hamiltonian. The reason for this is that at the crossover, both regimes with their excitations contribute, leading to larger bond dimensions. In contrast to the sharp peaks of the remnants of the QPTs (more precisely, ground state crossings), we expect the peaks in the QTT bond dimensions of the correlators at these thermal crossovers to be broader and their position to depend on temperature.
Furthermore, since Fourier transformation to Matsubara frequencies can be represented by a low-rank matrix product operator acting on the QTT representation of the imaginary-time correlator [19, 38], it can be expected that peaks in the bond dimension of the frequency representation match those of their imaginary times counterparts. Since at a thermal crossover or at a QPT multiple phases or regimes provide diverse contributions to correlation functions, more information has to be exchanged between the different length or time scales of the system leading to an increase in length or time scale entanglement. Therefore, the QTTD approach is not limited to specific two-point correlators, but can be utilized with arbitrary nonzero multi-point correlators due to the QTT bond dimension as a measure of system-inherent time scale entanglement. Moreover, the peaks in the bond dimensions of the QTT representation of multi-point correlators become more pronounced and stable with respect to various QTT cutoffs than in the two-point case (see the derivation in App. A). In App. B we also elaborate on how to treat low-accuracy data of correlators by varying the QTT cutoff to more precisely identify phase transitions and crossovers seen as stable peaks in $D_{\mathrm{max}}$.
We can summarize the QTTD procedure in the following way. i) Calculate an available correlator of the system. ii) Vary the QTT cutoffs to find stable peaks corresponding to possible phase transitions or crossovers. iii) Iterate on that procedure for different available correlators to verify the obtained maxima as universal features corresponding to the system-inherent rise in time scale entanglement associated with a phase transition or crossover. iv) Vary the temperature in the calculation of the correlators to disentangle ground state crossings from thermal crossovers. Moreover, since various physical properties, e.g. physical susceptibilities or entanglement, heavily depend on the dominant phase, features of these properties can also be visible in the maxima in the QTT bond dimension of the correlators.
IV Finite one-dimensional Hubbard models
Let us illustrate QTTD by attempting to find ground state crossings of a series of small benchmark systems. We start with the one dimensional Hubbard model on $N$ sites:
$\hat{H}$ $=t\sum_{\sigma,i=1}^{N}\left(\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{i+1,\sigma}+\hat{c}_{i+1,\sigma}^{\dagger}\hat{c}_{i,\sigma}\right)+t^{\prime}\sum_{\sigma,i=1}^{N}\left(\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{i+2,\sigma}+\hat{c}_{i+2,\sigma}^{\dagger}\hat{c}_{i,\sigma}\right)$
$+\sum_{i=1}^{N}\left(U\hat{n}_{i,\uparrow}\hat{n}_{i,\downarrow}-\mu(\hat{n}_{i,\uparrow}+\hat{n}_{i,\downarrow})\right),$ (5)
with the (next-)nearest neighbor hopping amplitude $t(t^{\prime})$, the local on-site Coulomb interaction $U$, the chemical potential $\mu$, the number operators $\hat{n}_{i,\sigma}=\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{i,\sigma}$ defined via the fermionic annihilation (creation) operators $\hat{c}_{i,\sigma}^{(\dagger)}$ with $\sigma=\uparrow,\downarrow$ and site index $i$. We use $t\equiv k_{B}\equiv 1$ to set energy and temperature units. In the following, we will study the cases of the Hubbard dimer and the four-site Hubbard ring with and without next-nearest neighbor hopping.
V Hubbard dimer
In the case of the Hubbard dimer [$N=2,t^{\prime}=0$ in Eq. (IV)], all 16 eigenstates and -values can be analytically calculated [39] and tracked, providing a first good testcase for the QTTD approach. Fig. 1 shows the bond dimension of the QTT representation of one- and two-particle Green’s functions as well as the corresponding spin susceptibility $\chi_{S}$ in the $\mu/U-U$-plane. In Fig. 1(a), the bond dimension of the local one-particle Green’s function $G_{11}^{\uparrow\uparrow}(\tau)=-\langle\mathcal{T}\hat{c}_{1,\uparrow}(\tau)\hat{c}_{1,\uparrow}^{\dagger}\rangle$ is shown at $\beta=50,\epsilon=2\cdot 10^{-6},R=12$. It can be seen that due to the small eigenspace (only 16 eigenvalues) and the cutoff in the QTT compression the bond dimension is rather small, because only few exponentials contribute. However, a peak in the bond dimension can be seen in the vicinity of the ground state crossover, indicated by the dashed lines. Here, the singlet ground state at half filling crosses with a doublet that has one (three) electrons at small (large) values of $\mu/U$. The picture becomes clearer when analyzing the bond dimension of the local two-particle Green’s function $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}(\tau_{1},\tau_{2},\tau_{3})=-\langle\mathcal{T}\hat{c}_{1,\uparrow}(\tau_{1})\hat{c}_{1,\uparrow}^{\dagger}(\tau_{2})\hat{c}_{1,\uparrow}(\tau_{3})\hat{c}_{1,\uparrow}^{\dagger}\rangle$ in Fig. 1 (b) with $\epsilon=10^{-14}$ at $\beta=50$ and $R=6$. Here, for three imaginary times more exponentials contribute giving rise to a smoother picture with larger bond dimensions. Thus, a sharp peak can be seen as a remnant of the QPT. Increasing temperature ($\beta=30$) in Fig. 1(d) leads to the expected smearing in the vicinity of the transition. In Fig. 1(c), the value of the spin susceptibility $\chi_{S}$ (see Appendix for definition) is shown, clearly distinguishing the two “phases”. In the Supplemental Material [40], we show that this analysis also applies to other correlators (including Matsubara frequency dependence) and temperatures, and also to crossings of the first excited states. We additionally demonstrate how to proceed with low-accuracy data by applying higher QTT cutoffs.
VI Four-site Hubbard ring
Next, let us consider the periodic 4-site Hubbard ring with nearest neighbor hopping only, i.e., $N=4,t^{\prime}=0$ in Eq. (IV). In Fig. 2(a), the bond dimension of the QTT representation of the two-particle Green’s function $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}(\tau_{1},\tau_{2},\tau_{3})$ is shown in the $\mu/U-U$-plane. Crossings of the ground state energy are indicated by red dots. For not too low values of $U$ it can be seen that the dots lie precisely on top of the peaks of the bond dimension. This becomes also visible in Fig. 2(e), where the slice for $U=4$ is shown and the remnants of the QPT (red dotted lines) can be seen as sharp peaks in $D_{\mathrm{max}}$. For lower values of $U$ the bond dimension peaks deviate from the crossings in the eigenspectrum of the Hamiltonian (red dashed lines) and become broader. This is because for small $U$ and $\beta=50$ excited states become thermally accessible. Consequently, when decreasing $T$, the region with large bond dimensions shifts to smaller $U$’s, see Fig. 2(f), and for $T\rightarrow 0$ again alienates with the eigenvalue crossings.
The spin susceptibility $\chi_{S}$ in Fig. 2(c) and the fidelity susceptibility $\chi_{F}$ (in Fig. 2(d) (see Appendix C for its definition) show a qualitatively similar behavior as the bond dimension: at large values of $U$ the ground state crossings are marked by a maximum, while for small values of $U$ both also show larger values, absent of an eigenvalue crossing. Further, when investigating entanglement in Fig. 2(b), also a similar structure is seen, where the white dashed line indicates the quantum Fisher information $F_{Q}=1$ (see Appendix for definition). Hence, the peaks in the bond dimension for low $U$ in Fig. 2(a) seem to be associated with some thermal crossover between different regimes. Further evidence of this thermal crossover of regimes seen by QTTD can be seen in Fig. 2(f), where the QTT bond dimension of $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}$ is shown for various temperatures. The dashed lines indicate where $F_{Q}=1$ and are located at the maxima of the different curves. Interestingly, also the derivative of the quantum Fisher information $\frac{dF_{Q}}{dU}$ (and other entanglement measures - see SM [40]) as well as that of the double occupancy show a maximum in close vicinity of this point. Additionally, also the maximum and inflection point of the spin susceptibility lie in proximity to this point and show a similar temperature behavior. We refer the reader to the SM [40] for further details and figures on all of the properties discussed.
Figure 2: Four-site Hubbard ring with nearest-neighbor hopping only: (a) $D_{\mathrm{max}}$ of the QTT for $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}(\tau_{1},\tau_{2},\tau_{3})$ at $\beta=50$, $\epsilon=10^{-14}$ and $R=6$ in comparison to (b) the quantum Fisher information $F_{Q}$, where the white dashed line indicates $F_{Q}=1$ , (c) spin susceptibility $\chi_{S}$ and (d) fidelity susceptibility. (e) and (f): $U=4$ and $\mu=U/2$ slice of (a), where red dashed lines in (e) indicate QPTs and dashed lines in (f) indicate $F_{Q}=1$. In (f), two additional $\beta$’s are shown.
VII Four-site Hubbard ring with next-nearest neighbor hopping
Next, as an indicator of time scale entanglement in the system, the bond dimension of the QTT representation of the two-particle correlator $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}$ of the periodic 4-site Hubbard ring with next-nearest neighbor hopping $t^{\prime}$ at $U=10$, $\beta=50$ is studied in the $t^{\prime}-\mu/U$-plane [41]. In contrast to the previous examples, we apply larger QTT cutoffs $\epsilon$. In Fig. 3(a), $D_{\mathrm{max}}$ of $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}(\tau_{1},\tau_{2},\tau_{3})$ is shown for $\epsilon=10^{-5}$. Still, the different ground state crossings (indicated by red dots) between s(inglet), d(oublet) and q(uartet) states lie precisely on top of the peaks of $D_{\mathrm{max}}$. This becomes even better visible in Figs. 3(b)-(c), where different $\mu/U$-slices are shown for a different QTT cutoff of $\epsilon=10^{-8}$. Here, the sharp peaks of the bond dimension are again precisely located at the QPTs demonstrating how well the proposed QTTD method works by producing stable results for different QTT cutoffs ($\epsilon=10^{-5}$ in (a) vs. $\epsilon=10^{-8}$ in (b)-(c)). This shows that the time scale entanglement diagnosed by the growth of the bond dimension $D_{\mathrm{max}}$ is system-inherent and visible across all studied correlators (see SM [40] additional plots). In App. B, a short recipe on how QTT cutoffs can be efficiently used, to deal with low-accuracy data is presented and in the SM [40], we comment on crossings that are correctly diagnosed by QTTD for which $\chi_{F}$ fails to achieve the same.
Figure 3: Four-site Hubbard ring with next nearest-neighbor hopping: Red dots (dotted lines) indicate crossings of s(inglet), d(oublet) and q(uartet) ground states. (a) $D_{\mathrm{max}}$ of QTT of $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}(\tau_{1},\tau_{2},\tau_{3})$ at $\beta=50$ with $\epsilon=10^{-5}$ and $R=6$. (b)-(c) $D_{\mathrm{max}}$ for various $\mu/U$-slices at $\epsilon=10^{-8}$.
VIII Single-impurity Anderson model
We examine the single-impurity Anderson model (SIAM), which features a magnetic impurity in a metallic bath and shows, among others, a crossover between local moment and Kondo regime. For numerical calculations, the continuous bath is discretized leading to the following Hamiltonian
$\hat{H}=$ $(\varepsilon_{0}-\mu)(\hat{n}_{\uparrow}+\hat{n}_{\downarrow})+U\hat{n}_{\uparrow}\hat{n}_{\downarrow}$
$+\sum_{\ell=1}^{3}E_{\ell}\hat{c}_{\ell}^{\dagger}\hat{c}_{\ell}+\sum_{\ell=1}^{3}(V_{\ell}\hat{d}_{\sigma}^{\dagger}\hat{c}_{\ell}+\mathrm{h.c.}).$ (6)
Here, $\varepsilon_{0}$ represents the impurity site energy, $\mu$ is the chemical potential, and $U$ is the on-site Coulomb interaction on the impurity site. The annihilation (creation) operator on the impurity site is denoted by $\hat{d}_{\sigma}^{({\dagger})}$, and the number operator is $\hat{n}_{\sigma}=\hat{d}_{\sigma}^{\dagger}\hat{d}_{\sigma}$. The density of states of the bath is assumed to be semicircular, and the parameters $V_{\ell}$ and $E_{\ell}$ denote the hybridization strength and energy levels of the discretized bath, respectively. The continuous bath is discretized into three levels, as detailed in the SM [40]. Figure 4(a) illustrates the bond dimension of the QTT representation of the two-particle Green’s function $G^{\uparrow\uparrow\uparrow\uparrow}_{\mathrm{imp}}(\tau_{1},\tau_{2},\tau_{3})$ at the impurity site for $\beta=100$, plotted against the model parameters $U$ and $V$, where $V$ denotes the hybridization strength in the original continuous model. In Fig. 4(b), the spin susceptibility $\chi_{S}$ is depicted. The red dots mark the Kondo temperature $T_{K}=V\sqrt{U}e^{-\pi U/8V^{2}}$ estimated via poor man’s scaling [42, 43], which delineates the boundary between the Kondo and local moment regimes, as also reflected in the spin susceptibility behavior. The bond dimension has a broad peak located in the vicinity of the Kondo temperature, diagnosing the thermally driven Kondo to local moments regime crossover.
Figure 4: (a) Bond dimension $D_{\mathrm{max}}$ of the two-particle Green’s function of the single-impurity Anderson model at $\beta=100$ in the $U$-$V$ plane, where the red dots indicate $T_{K}(U,V)=1/\beta$. (b) Spin susceptibility $\chi_{S}$ indicating the different regimes.
IX Conclusions
Since their recent invention, QTTs have primarily been applied as a compression tool in numerical calculations to overcome severe memory bottlenecks. In this work, we have presented an entirely novel approach on diagnosing phase transitions and crossovers with QTTs. We have demonstrated that this new quantics tensor train diagnostics approach is able to identify QPTs and thermal regime crossovers and distinguish between these two by studying various one-dimensional Hubbard rings and the SIAM. Moreover, by showing results for different correlators we have identified time scale entanglement to be system-inherent showing maxima at phase transitions and crossovers and, additionally, have pointed out that this method is also able to deal with low-accuracy data by varying the QTT cutoff. For the four-site ring with next-nearest neighbor hopping, QTTD correctly diagnosed ground state crossings that the fidelity susceptibility failed to capture. The strong point of the QTTD is that it can be used universally for zero and finite temperature and multiple correlation functions – it is not limited to ground state methods and does not require specific measures or susceptibilities. Further, QTTD offers a new perspective and deeper understanding on the inherent time and length scale entanglement at phase transitions.
Acknowledgements.
X Acknowledgements
We thank Samuel Badr, Jan von Delft, Herbert Eßl, Markus Frankenbach, Matthias Reitner, Marc Ritter and Nepomuk Ritz for insightful discussions. This work was funded in part by the Austrian Science Fund (FWF) projects through Grant DOI 10.55776/P36332, 10.55776/F86, 10.55776/I5868, and 10.55776/V1018. For open access purposes, the authors have applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission. Calculations have been partly performed on the Vienna Scientific Cluster (VSC). H.S. was supported by JSPS KAKENHI Grants No. 21H01041, No. 21H01003, and No. 23H03817 as well as JST FOREST Grant No. JPMJFR2232, Japan. H.I. and H.S. were supported by JST PRESTO Grant No. JPMJPR2012, Japan.
Appendix A Appendix A: QTTD and multi-point correlators
Let us consider an arbitrary n-point correlator function $\mathcal{G}(\bm{\tau})=-\langle\mathcal{T}\hat{O}_{1}(\tau_{1})\hat{O}_{2}(\tau_{2})...\hat{O}_{n}(0)\rangle$ with $\bm{\tau}=(\tau_{1},\tau_{2},...,\tau_{n-1}),\tau_{i}\in[0,\beta)$ and fermionic or bosonic operators $\hat{O}_{i}$. We express the correlator in terms of its Lehmann representation
$\mathcal{G}(\bm{\tau})$ $=-\frac{1}{Z}\sum_{m_{1}}e^{-\beta\epsilon_{m_{1}}}\sum_{m_{2}...m_{n}}e^{-\tau_{1}(\epsilon_{m_{2}}-\epsilon_{m_{1}})}e^{-\tau_{2}(\epsilon_{m_{3}}-\epsilon_{m_{2}})}...e^{-\tau_{n-1}(\epsilon_{m_{n}}-\epsilon_{m_{n-1}})}$
×$\times\langle m_{1}|\hat{O}_{1}|m_{2}\rangle\langle m_{2}|\hat{O}_{2}|m_{3}\rangle...\langle,m_{n}|\hat{O}_{n}|m_{1}\rangle$ (7)
with the eigenstates $|m_{i}\rangle$ of the Hamiltonian and we set the ground state energy $\epsilon_{\mathrm{GS}}=0$ in the following. Then, in the case of temperature approaching zero ($\beta\rightarrow\infty$) only summands including the ground state $|\mathrm{GS}\rangle$ ($\exists i,i=1,...,n:|m_{i}\rangle=|\mathrm{GS}\rangle$) will give non-zero contributions, because all other contributions are exponentially suppressed. If the temperature is finite, but very low, also the other summands will contribute, however they will still be heavily suppressed. Therefore, the correlator will still be primarily dominated by terms containing the ground state. E.g. in the case of a three-point correlator, the following terms are dominating
$|m_{1}\rangle=|\mathrm{GS}\rangle:$ (8a)
$\quad e^{-\tau_{1}\epsilon_{m_{2}}}e^{-\tau_{2}(\epsilon_{m_{3}}-\epsilon_{m_{2}})}\langle\mathrm{GS}|\hat{O}_{1}|m_{2}\rangle\langle m_{2}|\hat{O}_{2}|m_{3}\rangle\langle m_{3}|\hat{O}_{3}|\mathrm{GS}\rangle,$
$|m_{2}\rangle=|\mathrm{GS}\rangle:$ (8b)
$\quad e^{-(\beta-\tau_{1})\epsilon_{m_{1}}}e^{-\tau_{2}\epsilon_{m_{3}}}\langle m_{1}|\hat{O}_{1}|\mathrm{GS}\rangle\langle\mathrm{GS}|\hat{O}_{2}|m_{3}\rangle\langle m_{3}|\hat{O}_{3}|m_{1}\rangle,$
$|m_{3}\rangle=|\mathrm{GS}\rangle:$ (8c)
$\quad e^{-(\beta-\tau_{1})\epsilon_{m_{1}}}e^{-(\tau_{1}-\tau_{2})\epsilon_{m_{2}}}\langle m_{1}|\hat{O}_{1}|m_{2}\rangle\langle m_{2}|\hat{O}_{2}|\mathrm{GS}\rangle\langle\mathrm{GS}|\hat{O}_{3}|m_{1}\rangle,$
where we already set $\epsilon_{\mathrm{GS}}=0$. Similar to the analysis of Eq. (III), the first summand dominates in the case of small $\tau_{1},\tau_{2}$, the second one in the case of large $\tau_{1}$ (close to $\beta$) and small $\tau_{2}$ and the third for large $\tau_{1},\tau_{2}$. In contrast to the case of the two-point correlator, it can be seen that apart from the dependence on two instead of one imaginary times, in general, more exponentials will contribute in the three-point correlator. Since the QTT representation of an exponential has a bond dimension of one and the bond dimension of the sum of two QTTs is bound by the sum of the bond dimensions of the individual QTTs, we expect three-point correlators to have larger QTT bond dimensions than two-point correlators due to additional imaginary time dependence and more dominating contributions in the sum. Therefore, higher-point correlators are also expected to show more pronounced peaks at phase transitions and crossovers, which is confirmed by the results in the paper. The reasoning that phase transitions and crossovers are connected to maxima in the bond dimensions of correlators follows the same arguments as in the main text. Let us emphasize that this derivation is valid for arbitrary (non-zero) correlators. Since Fourier transformation can be represented by a low-rank matrix product operator applied to the correlator QTT, imaginary frequency counterparts to the imaginary times correlator are expected to display the same maxima in the bond dimension, which is supported by the results in the Supplemental Material. Therefore, arbitrary correlators that are available from computations can be used in QTTD to diagnose phase transitions and crossovers in systems. Furthermore, since the bond dimension is a measure of time scale entanglement, we argue that time scale entanglement is primarily a system-inherent property, because arbitrary correlators share the same bond dimension features.
Appendix B Appendix B: Low-accuracy data and diagnostics
In contrast to many conventional methods of identifying phase transitions and crossovers, the choice of the QTT cutoff allows for flexibility in dealing with low-accuracy data. This can be understood in the following way. Let us consider the case of noise in the evaluation of the correlators (e.g. in our case by truncation of the eigenspace). Then the bond dimension of a QTT compression with a cutoff or tolerance below the noise level will suffer from artifacts due to the inclusion of the noise in the compressed QTT. However, these artifacts will be highly dependent on the chosen cutoff. If the QTT cutoff is increased above the noise level, the artifacts will vanish and only peaks in the QTT bond dimensions associated with phase transitions and crossovers will remain. These peaks are then very stable for a broad range of different cutoffs above the error level clearly indicating the underlying maxima in the length and time scale entanglement of the system. Therefore, it is possible to not only identify low-accuracy input data, but to overcome its limitations by truncating the noise in the QTT compression. Further evidence and support of this property is provided in the SM [40].
Appendix C Appendix C: Fidelity susceptibility
Following Ref. 44, we define the fidelity susceptibility [45, 46, 47, 48, 49] at finite temperatures as
[…]
with
$\hat{H}(\lambda)=\hat{H}_{0}+\lambda\hat{H}_{1},$ (10)
where $\lambda$ is the “driving parameter”, $\hat{H}_{0}$ the non-interacting part and $\hat{H}_{1}$ the interacting part of the Hamiltonian. The fidelity susceptibility describes how susceptible the overlap of neighboring ground states (or density matrices in the case of finite temperatures) in parameter space is to small (driving) parameter changes. Similarly to QTTD, the fidelity susceptibility is expected to show maxima at phase transitions. For our considerations, we will consider the Coulomb interaction $U$ to be the driving parameter of the phase transitions.
Appendix D Appendix D: Diagnostics tool
In the main text, this technique was coined “quantics tensor train diagnostics”. Here, let us prove the “diagnostics” character of the approach for gapped systems (insulating or finite systems) at $T=0$. For this reason, we will revisit Eq. (III). Then, as discussed, the number of exponentials in the sum depends on the matrix elements $\langle\mathrm{GS}_{i}|\hat{A}|m\rangle\langle m|\hat{B}|\mathrm{GS}_{i}\rangle$ and $\langle m|\hat{A}|\mathrm{GS}_{i}\rangle\langle\mathrm{GS}_{i}|\hat{B}|m\rangle$ being nonzero. However, within one “phase” (ground state regime) in parameter space, the number of nonzero terms is fixed because gapped systems are considered without continuous ground state changes. This results in the same number of exponentials contributing to the correlation function, which we denote by $\mathcal{N}_{1}$. Although, individual contributions might become increasingly suppressed within one ground state regime, they will not entirely vanish leading to a fixed bond dimension matching the number of exponentials in the exact analytical construction of the corresponding QTT. Considering a neighboring but differing phase, the same argument is valid, resulting in the contribution of $\mathcal{N}_{2}$ exponentials, leading to a bond dimension of $\mathcal{N}_{2}$. Then a maximum in the bond dimension bound by $\mathcal{N}_{1}+\mathcal{N}_{2}$ from above is observed precisely at the crossing of the two ground states, where the different exponentials from both ground states contribute. If the two ground states do not have any matching states $|m\rangle$ that lead to nonzero contributions, the bond dimension in the exact analytical construction at the boundary will be precisely $\mathcal{N}_{1}+\mathcal{N}_{2}$. Let us mention that in both phases it might also be possible to see small dips in $D_{\mathrm{max}}$, if two excited states $|m\rangle$ that both lead to nonzero contributions cross, because the contributing exponentials of both states will match each other. Hence, at $T=0$ a maximum in the bond dimension of the exact analytical QTT construction of an arbitrary correlator corresponds to a crossing of two ground state regimes. However, at finite temperatures, all states will contribute to the correlator, which would lead to a constant bond dimension in the entire parameter space in the exact analytical construction. Since at low temperatures most of the contributions will be heavily suppressed, they will be truncated in a numerical QTT construction of the correlator, restoring the diagnostics power of the presented method. However, truncation can come at the price of varying bond dimensions within one ground state regime, which can lead to broader and smaller maxima, also within one ground state regime. Since the system-inherent amount of length and time scale entanglement is the driving force of the maximum of the bond dimension at a ground state crossing, by studying various correlators and varying the QTT cutoffs and also the temperature, the diagnostics power can be fully restored, where clear stable sharp peaks can be associated with ground state crossings, while broader temperature-dependent peaks are corresponding to thermal crossings.
Appendix E Appendix E: Spin and charge susceptibility
The local spin susceptibility $\chi_{S}$ is defined in the following way [43]
[…]
where $\hat{S}_{z}=(\hat{n}_{\uparrow}-\hat{n}_{\downarrow})/2$ describes the local magnetization on a single site. Similarly, the local charge susceptibility is defined as follows
[…]
with $\hat{n}=(\hat{n}_{\uparrow}+\hat{n}_{\downarrow})/2$.
Appendix F Appendix F: Entanglement measures
The quantum Fisher information (QFI) [50]
$F_{Q}\equiv\int_{0}^{\infty}\tanh\left(\frac{\omega\beta}{2}\right)\textrm{Im}\chi(\omega)$ (13)
serves as an experimentally accessible measure of multipartite entanglement. A value $F_{Q}\geq m$ detects $m+1$-partite entanglement. A peak in the derivative of the QFI hints at a phase transition [51].
The two-site mutual information is a measure of total correlations between two lattice sites based on the von Neumann entropy[52, 53]
$I_{ij}\equiv\textrm{tr}[\rho_{ij}\ln\rho_{ij}]-2\textrm{tr}[\rho_{i}\ln\rho_{i}],$ (14)
where $\rho_{ij}$ is the two-site reduced density matrix and $\rho_{i}=\rho_{j}=\textrm{tr}_{i}[\rho_{ij}]$ [54, 55, 56]. This can be generalized to the Rényi mutual information [57, 58]
$I_{R}\equiv\ln\left[\textrm{tr}\rho_{ij}^{2}\right]-2\ln\left[\textrm{tr}\rho_{i}^{2}\right].$ (15)
The partial transpose $\rho_{ij}^{T_{j}}$ of the two-site reduced density matrix provides another entanglement measure [59, 60]. If the density matrix is separable, then $\rho_{ij}^{T_{j}}=\rho_{i}\otimes\rho_{j}^{T}$ is a physical density matrix that is positive semi-definite. Vice versa, negative eigenvalues $\lambda$ of $\rho_{ij}^{T_{j}}$ indicate entanglement and the negativity is defined as
$N\equiv\sum_{\lambda<0}|\lambda|.$ (16)
In a phase transition, the density matrix undergoes a rapid change. We illustrate this with a trivial example, in a zero temperature transition from a paramagnetic to a ferromagnetic state, all antiferromagnetic contributions vanish and the density matrix has a discontinuous jump. At higher temperatures this will be less sharp as excited states may still contribute. Similarly, in a crossover, a smooth change between the density matrices of the two regimes is expected. Since the two-site mutual information and the negativity are directly computed from the density matrix, such a change shows as an extrema in the derivatives of these measures as well [61, 62, 63].

### v2

@@
 Abstract
-We propose a new method based on the compression of arbitrary multi-point correlators into quantics tensor trains and studying its so called bond dimension as an indicator of time scale entanglement in the system.
-This quantics tensor train diagnostics approach allows us to i) identify quantum phase transitions (ground state crossings) and thermal crossovers, ii) distinguish between these two by allowing flexibility through iii) using arbitrary available correlators and iv) the possibility to deal with low-accuracy input data through varying the quantics tensor train cutoffs.
-We verify the validity of the approach by studying the phase diagram of the Hubbard dimer, the four-site Hubbard ring with and without next-nearest neighbor hopping and the single-impurity Anderson model.
-We find time scale entanglement to be a system-inherent property becoming maximal at phase transitions and crossover, which translates to peaks in the quantics tensor train bond dimension of correlator functions.
-Additionally, the connection of the method to other physical quantities, such as various entanglement measures and susceptibilities, is shown.
+Spatial entanglement of wave functions has matured into an enthralling and very active research area.
+Here, we unearth a completely different kind of entanglement, the entanglement between different time scales.
+This is feasible through quantics tensor train diagnostics (QTTD), wherein the bond dimension for an $n$-particle correlation function allows diagnosing the temporal entanglement.
+As examples, we study time-scale entanglement of the Hubbard dimer, the four-site Hubbard ring with and without next-nearest neighbor hopping and the single-impurity Anderson model.
+Besides introducing the QTTD method, our major finding is that the time-scale entanglement is generically maximal at phase transitions and crossovers.
+This is independent of the correlation function studied.
+Thus, QTTD is a universal tool for detecting quantum phase transitions, ground state crossings in finite systems, and thermal crossovers.
 preprint: APS/123-QED
@@
 Phase transitions are among the most intriguing phenomena in physics.
-Yet, their numerical treatment, particularly in systems with nontrivial electronic behavior, can be challenging, especially when exotic phases emerge [1, 2, 3].
-In particular, the full wave function is not available for many systems, and using susceptibilities to diagnose transitions can be brittle, since we need to know a priori where to look, i.e., we must make sure to compute the response associated with the transition in question or at least one that is not insensitive to the transition.
+Yet, their numerical treatment and detection can be challenging, especially when exotic phases emerge [1, 2, 3].
+In extended many-body systems, the full wave function is usually not available, and using susceptibilities to diagnose transitions can be brittle, since we need to know a priori where to look, i.e., we must make sure to compute the response associated with the transition in question or at least one that is not insensitive to the transition.
 In general, we can expect some but a much weaker signal in the susceptibilities that are not the one associated with the symmetry breaking for a second order phase transition.
-A first order phase transitions, on the other hand, will not be signaled at all in the susceptibilities prior to the phase transition.
-Tensor trains, used originally as matrix product states (MPSs) for the ground state of spin systems [4, 5], have been elevated in the last years to a general compression ansatz [6, 7, 8].
-By combining tensor trains with the quantics representation – a set of auxiliary qubits for exponentially different time or length scales – one obtains a compression ansatz for arbitrary functions in space and time [9, 10, 11] called a quantics tensor train (QTT).
-QTTs were instrumental, among others, in computational breakthroughs in modeling of turbulence [12, 13, 14, 15, 16], plasma physics [17], and quantum chemistry [18].
-Some of the authors of this Letter have recently extended the use of QTTs for modeling electronic correlation functions [19], and found that it unifies prior pole [20, 21], basis [22, 23, 24, 25], and mesh refinement [26] ansätze into a single framework – one where doubling the size of the system amounts to adding a qubit, a mere linear increase.
-When combined with a way to sample QTTs [27, 28], this framework allows the solution of diagrammatic theories for systems which were previously inaccessible [29].
-Moreover, full $n$-particle correlators can be compressed and stored.
-The compression power of tensor trains depends on the connection strength between the qubits and thus between the length or time scales, encoded in the so-called bond dimension $D_{\mathrm{max}}$.
-For physical qubits used in MPSs for spin chains, this bond dimension has physical significance: it is related to the entanglement between two parts of the system [30]; for auxiliary qubits used in QTTs for correlation functions, it has until now only been used as a performance metric.
-Figure 1: QTT bond dimension $D_{\mathrm{max}}$ of different imaginary-time correlation functions for the Hubbard dimer as a function of electronic repulsion $U$ and chemical potential $\mu/U$: (a) $D_{\mathrm{max}}$ for the one-particle Green’s function $G_{11}^{\uparrow\uparrow}$ at temperature $\beta^{-1}=1/50$, and (b,d) for the two-particle Green’s function $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}$ at $\beta^{-1}=1/50$ and $1/30$, respectively.
-For comparison, the value of the static, local spin susceptibility $\chi_{s}$ at $\beta^{-1}=1/50$ is shown in panel (c).
-Red dashed lines mark crossings of the ground state of the model.
-These crossings can be diagnosed as a sharp maximum in $D_{\mathrm{max}}$.
-In this Letter, we argue that the bond dimension of correlation functions has physical significance: it becomes maximal at crossovers and phase transitions [red dots in Fig.
-1].
-These maxima appear in both single- [Fig.
-1 (a)] and two-particle propagators [Fig.
-1 (b,c)], without the need to identify the specific susceptibility associated with the transition.
+A first order phase transition, on the other hand, will not be signaled at all in the susceptibilities prior to the phase transition.
+Similarly, entanglement is one of the most enigmatic quantum phenomena.
+It is equally challenging to calculate in true many-electron systems where many electrons contribute and are entangled [4, 5, 6, 7, 8, 9, 10, 11, 12, 13].
+In (quasi-) one-dimensional spin systems, matrix product states (MPS) [14, 15] have unlocked much of the theoretical and numerical analysis of entanglement [16, 17] and its growth [18, 19, 20].
+MPS add to the physical space an auxiliary, latent space, whose dimension (“bond dimension”) indicates the strength of spatial entanglement [21] [22].
+Recently, it was shown that the machinery of MPS can be reused to compactify arbitrary functions in space and time [23, 24, 25, 26, 27, 28].
+Instead of a physical spin at some site, each tensor represents a different time/length scale of the function, so as we step along the train, we zoom in and out rather than move left or right.
+This ansatz was termed quantics tensor train (QTT).
+It was instrumental in breaking computational barriers in modeling turbulence [29, 30, 31, 32, 33], plasmas [34], quantum chemistry [35], and – most relevant to this Letter – electronic correlation functions.
+In this latter case, some of us showed how to sample [36, 37], store [38], and perform calculations [39] with electronic response functions which were previously inaccessible.
+QTTs by construction also feature a latent space and an associated bond dimension, which has hitherto been chiefly used as a measure of performance.
+Figure 1: Entanglement of exponentially different time scales for the Hubbard dimer at $T=0.02$: (a) Local spin susceptibility $\chi_{s}$ indicating different ground states filled with one, two and three electrons; (b) QTT maximum bond dimension $D_{\max}:=\max_{\ell}D_{\ell}$ of the imaginary-time four-point Green’s function $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}$ as a function of interaction strength $U$ and chemical potential $\mu/U$, where red-dashed lines mark crossings of the ground state; and (c) QTT bond dimension $D_{\ell}$ between timescale $2^{-\ell}$ and $2^{-\ell-1}$ for different values of $\mu/U$ at $U=8$, marked by the same symbol and color in (a) and (b).
+In this Letter, we reason by analogy with MPS, where the bond dimension gauges spatial entanglement: we propose the QTT bond dimension of electronic functions as a physical measure, namely as the strength of time and length scale entanglement.
+For the example of the Hubbard dimer, we plot this over different time scales in Fig.
+1(c).
+Furthermore, we observe that the maximal bond dimension $D_{\mathrm{max}}$ shows a maximum at crossovers and phase transitions [red dots in Fig.
+1(b)].
+These maxima appear in both single- and two-particle propagators [Fig.
+1 (b)], without the need to identify the specific susceptibility associated with the transition.
 We call this program of detecting phase transitions quantics tensor train diagnostics (QTTD).
 II Quantics tensor trains
-In the QTT approach [6, 19] each variable is represented through a set of binary numbers or “quantics” corresponding to different length or time scales.
+In QTTs  [23, 38] each variable is represented through a set of binary numbers or “quantics” corresponding to different length or time scales.
 The resulting tensor is then factorized into a tensor train (TT) at each scale.
-To illustrate the approach, let us consider a function $f(\tau)$ of a discretized variable $\tau$, with $\tau\in\{0,\ldots,M-1\}$ on $M=2^{R}$ grid points.
-In the quantics approach, $\tau$ is expressed in its binary representation $\tau=(\sigma_{1}\sigma_{2}\dots\sigma_{R})_{2}=\sum_{\ell=1}^{R}2^{R-\ell}\sigma_{\ell},$ with $\sigma_{\ell}\in\{0,1\}\,$ leading to $f(\tau)$ being seen as a $2\times 2\times...\times 2$ ($R$ times) tensor $F_{\sigma_{1},\ldots,\sigma_{R}}$ instead.
+For illustration, consider a function $f(\tau)$ of a discretized variable $\tau$, with $\tau\in\{0,\ldots,M-1\}$ on $M=2^{R}$ grid points.
+In quantics, $\tau$ is expressed in binary representation $\tau=(\sigma_{1}\sigma_{2}\dots\sigma_{R})_{2}=\sum_{\ell=1}^{R}2^{R-\ell}\sigma_{\ell},$ with $\sigma_{\ell}\in\{0,1\}\,$ leading to $f(\tau)$ being seen as a $2\times 2\times...\times 2$ ($R$ times) tensor $F_{\sigma_{1},\ldots,\sigma_{R}}$ instead.
 Now each tensor index $\sigma_{\ell}$ corresponds to an exponentially distinct length or time scale of the system.
@@
 $F_{\sigma_{1}\ldots\sigma_{R}}\approx\tilde{F}_{\sigma_{1}\ldots\sigma_{R}}=\!\sum_{\alpha_{1}=1}^{D_{1}}\!\ldots\!\sum_{\alpha_{R-1}=1}^{D_{R-1}}[M_{1}]^{\sigma_{1}}_{1\alpha_{1}}[M_{2}]^{\sigma_{2}}_{\alpha_{1}\alpha_{2}}\cdots[M_{R}]^{\sigma_{R}}_{\alpha_{R-1}1}$ (1)
-can be achieved by means of singular value decomposition (SVD) [19] or using the tensor cross interpolation (TCI) algorithm [8, 28, 27].
+can be achieved with singular value decomposition (SVD) [38] or using tensor cross interpolation (TCI) [25, 37, 36].
 In Eq.
@@
 [Uncaptioned image] (2)
-in tensor network form [5].
-We define the (maximum) bond dimension of the QTT as $D_{\mathrm{max}}=\textrm{max}_{\ell}(D_{\ell})$.
-Generally, the bond dimensions $D_{\ell}$ characterize the amount of “length or time scale entanglement” in the system are truncated in the TT factorization such that a specified error cutoff or tolerance $\epsilon$ is satisfied.
-In the case of SVD, usually the squared Frobenius norm $\epsilon=||F-\tilde{F}||^{2}_{\mathrm{F}}/||F||^{2}_{\mathrm{F}}$ is taken as a measure of the error of the approximation in Eq.
-(1).
-Many problems in many-body physics have been shown to use quantities that are strongly compressible with the QTT approach, which is possible if the bond dimension is sufficiently small [19, 31, 32, 33, 28, 29, 34, 35, 36, 37].
+[…]
 III Quantics tensor trains diagnostics
-Let us start from a general imaginary time two-point correlator $G_{AB}(\tau)=-\langle\hat{\mathcal{T}}\hat{A}(\tau)\hat{B}\rangle$ with the fermionic operators $\hat{A},\hat{B}$ and the time ordering operator $\hat{\mathcal{T}}$.
+Let us start from a general imaginary time two-point correlator $G_{AB}(\tau)=-\langle\hat{\mathcal{T}}\hat{A}(\tau)\hat{B}\rangle$ with fermionic operators $\hat{A},\hat{B}$ and the time ordering operator $\hat{\mathcal{T}}$.
 Then, $G_{AB}(\tau)$ can be expressed in its Lehmann representation:
@@
 where $1/\beta$ is temperature, $\tau\in[0,\beta)$ is imaginary time, $Z$ is the partition function, $\hat{H}$ is the Hamiltonian including the chemical potential, and $\hat{H}|n\rangle=\epsilon_{n}|n\rangle$ defines its eigenenergies and eigenstates.
-As can be seen, $G_{AB}(\tau)$ is the sum of exponentials.
-A single exponential has a QTT representation of bond dimension one, since it is a product of exponentials of the binary variables $\sigma_{l}$ at each scale: $e^{\tau\epsilon}=\prod_{\ell=1}^{R}e^{\epsilon\beta 2^{-\ell}\sigma_{\ell}}$.
+Clearly, $G_{AB}(\tau)$ is the sum of exponentials.
+An exponential has a QTT representation of bond dimension one, since it is a product of exponentials of the binary variables $\sigma_{l}$ at each scale: $e^{\tau\epsilon}=\prod_{\ell=1}^{R}e^{\epsilon\beta 2^{-\ell}\sigma_{\ell}}$.
 The bond dimension of the exact QTT representation of $G_{AB}$ is, thus, bound by the number of nonzero elements in the sum.
@@
 We set $\epsilon_{\mathrm{GS}}=0$; and the first (second) term contributes predominately for small (large) values of $\tau$.
-First, let us focus on the case of a single non-degenerate ground state.
-Then, the bond dimension of the exact QTT representation is bound by the number of nonzero terms in the sum, so at maximum by $M-1$ (if all matrix elements are nonzero) for $M$ eigenstates $m$.
+For a single non-degenerate ground state, the bond dimension of the exact QTT representation is bounded by the number of nonzero terms in the sum, maximally by M1M-1 for $M$ eigenstates $m$.
 Next, consider a ground state crossing of the two lowest states $|\mathrm{GS}_{1}\rangle$ and $|\mathrm{GS}_{2}\rangle$ ($N=2$).
-Then the bond dimension of the exact QTT is at most $22M-$ 2 since twice as many exponentials can contribute.
-In practice, it may be lower as only some of the states $|m\rangle$ will give nonzero matrix elements.
+Then the bond dimension of the exact QTT is at most 2M22M-2 since twice as many exponentials can contribute.
+In practice, it may be lower as fewer states $|m\rangle$ will give nonzero matrix elements.
 Still, at the ground state crossing the sum consists of the exponentials and matrix elements associated with $|\mathrm{GS}_{1}\rangle$ and $|\mathrm{GS}_{2}\rangle$ leading to larger bond dimensions of $G_{AB}(\tau)$ than before or after the crossing.
-Therefore, at zero temperature, a quantum phase transition (QPT) or a ground state crossing of the Hamiltonian can be determined by monitoring the bond dimension of the QTT representation.
-We will call this approach quantics tensor train diagnostics (QTTD), where the diagnostics power of the approach is further examined in Appendix D.
+Therefore, at zero temperature, a quantum phase transition (QPT) or a ground state crossing can be determined by monitoring the bond dimension of the QTT representation.
+We call this approach quantics tensor train diagnostics (QTTD), where the diagnostics power of the approach is further examined in Appendix D.
 Similarly, ground state crossings of degenerate states (e.g.
-a doublet with a triplet state) will lead to maxima of the bond dimension in parameter space.
-Let us now consider finite, but low temperatures.
+a doublet with a triplet state) will lead to maxima of $D_{\mathrm{max}}$ in parameter space.
+Consider finite, but low temperatures.
 Then, also excited states $|n\rangle$ contribute to $G_{AB}(\tau)$, suppressed by $e^{-\beta\epsilon_{n}}$.
-However, if we compress the two-point correlator into a QTT, most of these contributions will be truncated, if they are below a certain specified cutoff in the QTT construction.
-Hence, we will still see a peak of the maximum bond dimension in the vicinity of the corresponding QPT (at $T=0$), which will become less sharp with increasing temperature which smears out this transition.
-The QTTD approach can also be used to identify thermal crossovers of different regimes, such as the Kondo effect, that cannot be identified by merely looking at the lowest energy states in the ($T$-independent) eigenspectrum of the Hamiltonian.
-The reason for this is that at the crossover, both regimes with their excitations contribute, leading to larger bond dimensions.
-In contrast to the sharp peaks of the remnants of the QPTs (more precisely, ground state crossings), we expect the peaks in the QTT bond dimensions of the correlators at these thermal crossovers to be broader and their position to depend on temperature.
-Furthermore, since Fourier transformation to Matsubara frequencies can be represented by a low-rank matrix product operator acting on the QTT representation of the imaginary-time correlator [19, 38], it can be expected that peaks in the bond dimension of the frequency representation match those of their imaginary times counterparts.
-Since at a thermal crossover or at a QPT multiple phases or regimes provide diverse contributions to correlation functions, more information has to be exchanged between the different length or time scales of the system leading to an increase in length or time scale entanglement.
-Therefore, the QTTD approach is not limited to specific two-point correlators, but can be utilized with arbitrary nonzero multi-point correlators due to the QTT bond dimension as a measure of system-inherent time scale entanglement.
-Moreover, the peaks in the bond dimensions of the QTT representation of multi-point correlators become more pronounced and stable with respect to various QTT cutoffs than in the two-point case (see the derivation in App.
+Most of these contributions will be truncated, if they are below a certain specified cutoff in the QTT construction of the two-point correlator.
+Hence, we will still see a peak of $D_{\mathrm{max}}$ in the vicinity of the corresponding QPT (at $T=0$), that will be smeared out with increasing temperature.
+QTTD can also identify thermal crossovers of different regimes, such as the Kondo effect, that cannot be identified by merely looking at the lowest energy states in the ($T$-independent) eigenspectrum of the Hamiltonian.
+At the crossover, both regimes with their excitations contribute, leading to larger bond dimensions.
+In contrast to the sharp peaks of the remnants of QPTs (more precisely, ground state crossings), we expect the peaks of $D_{\mathrm{max}}$ of the correlators at thermal crossovers to be broader and their position to depend on temperature.
+Since at thermal crossovers or at QPTs multiple phases or regimes provide diverse contributions to correlation functions, more information has to be exchanged between different length or time scales of the system leading to an increase in length or time scale entanglement.
+Therefore, QTTD is not limited to specific two-point correlators, but can be utilized with arbitrary nonzero multi-point correlators due to the QTT bond dimension as a measure of system-inherent time scale entanglement.
+Moreover, peaks in $D_{\mathrm{max}}$ of the QTT representation of multi-point correlators become more pronounced and stable with respect to various QTT cutoffs than in the two-point case (see derivation in App.
 A).
 In App.
-B we also elaborate on how to treat low-accuracy data of correlators by varying the QTT cutoff to more precisely identify phase transitions and crossovers seen as stable peaks in $D_{\mathrm{max}}$.
+B, we also elaborate on how to treat low-accuracy data of correlators by varying the QTT cutoff to more precisely identify phase transitions and crossovers seen as stable peaks in $D_{\mathrm{max}}$.
 We can summarize the QTTD procedure in the following way.
@@
 ii) Vary the QTT cutoffs to find stable peaks corresponding to possible phase transitions or crossovers.
-iii) Iterate on that procedure for different available correlators to verify the obtained maxima as universal features corresponding to the system-inherent rise in time scale entanglement associated with a phase transition or crossover.
+iii) Iterate on this procedure for different available correlators to verify the obtained maxima as universal features corresponding to the system-inherent rise in time scale entanglement associated with a phase transition or crossover.
 iv) Vary the temperature in the calculation of the correlators to disentangle ground state crossings from thermal crossovers.
@@
 IV Finite one-dimensional Hubbard models
-Let us illustrate QTTD by attempting to find ground state crossings of a series of small benchmark systems. We start with the one dimensional Hubbard model on $N$ sites:
-$\hat{H}$ $=t\sum_{\sigma,i=1}^{N}\left(\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{i+1,\sigma}+\hat{c}_{i+1,\sigma}^{\dagger}\hat{c}_{i,\sigma}\right)+t^{\prime}\sum_{\sigma,i=1}^{N}\left(\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{i+2,\sigma}+\hat{c}_{i+2,\sigma}^{\dagger}\hat{c}_{i,\sigma}\right)$
-$+\sum_{i=1}^{N}\left(U\hat{n}_{i,\uparrow}\hat{n}_{i,\downarrow}-\mu(\hat{n}_{i,\uparrow}+\hat{n}_{i,\downarrow})\right),$ (5)
-with the (next-)nearest neighbor hopping amplitude $t(t^{\prime})$, the local on-site Coulomb interaction $U$, the chemical potential $\mu$, the number operators $\hat{n}_{i,\sigma}=\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{i,\sigma}$ defined via the fermionic annihilation (creation) operators $\hat{c}_{i,\sigma}^{(\dagger)}$ with $\sigma=\uparrow,\downarrow$ and site index $i$.
+We start with the one dimensional Hubbard model on $N$ sites:
+$\hat{H}=-\sum_{\sigma}\sum_{i,j=1}^{N}t_{i-j}\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{j,\sigma}+\sum_{i=1}^{N}\left(U\hat{n}_{i,\uparrow}\hat{n}_{i,\downarrow}-\mu(\hat{n}_{i,\uparrow}+\hat{n}_{i,\downarrow})\right),$ (5)
+where $t_{\pm 1}\equiv t$ and $t_{\pm 2}\equiv t^{\prime}$ are the nearest and next-nearest hopping amplitude, respectively, all other hopping amplitudes are set to zero, the local on-site Coulomb interaction is $U$, the chemical potential is $\mu$, and the number operators $\hat{n}_{i,\sigma}=\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{i,\sigma}$ are defined via the fermionic annihilation (creation) operators $\hat{c}_{i,\sigma}^{(\dagger)}$ with $\sigma=\uparrow,\downarrow$ and site index $i$.
 We use $t\equiv k_{B}\equiv 1$ to set energy and temperature units.
@@
 V Hubbard dimer
-In the case of the Hubbard dimer [$N=2,t^{\prime}=0$ in Eq.
-(IV)], all 16 eigenstates and -values can be analytically calculated [39] and tracked, providing a first good testcase for the QTTD approach.
+For Hubbard dimer $N=2,t^{\prime}=0$ in Eq.
+(5)], all 16 eigenstates can be analytically calculated [48], providing a first good testcase for QTTD.
 Fig.
-1 shows the bond dimension of the QTT representation of one- and two-particle Green’s functions as well as the corresponding spin susceptibility $\chi_{S}$ in the $\mu/U-U$-plane.
+1(b) shows $D_{\mathrm{max}}$ of the local two-particle Green’s function $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}(\tau_{1},\tau_{2},\tau_{3})=-\langle\mathcal{T}\hat{c}_{1,\uparrow}(\tau_{1})\hat{c}_{1,\uparrow}^{\dagger}(\tau_{2})\hat{c}_{1,\uparrow}(\tau_{3})\hat{c}_{1,\uparrow}^{\dagger}\rangle$ with $\epsilon=10^{-14}$ at $\beta=50$ and $R=6$.
+A sharp peak in $D_{\mathrm{max}}$ can be seen in the vicinity of the ground state crossover, indicated by dashed lines.
+Here, the singlet ground state crosses with a doublet that has one (three) electrons at small (large) values of $\mu/U$.
 In Fig.
-1(a), the bond dimension of the local one-particle Green’s function $G_{11}^{\uparrow\uparrow}(\tau)=-\langle\mathcal{T}\hat{c}_{1,\uparrow}(\tau)\hat{c}_{1,\uparrow}^{\dagger}\rangle$ is shown at $\beta=50,\epsilon=2\cdot 10^{-6},R=12$.
-It can be seen that due to the small eigenspace (only 16 eigenvalues) and the cutoff in the QTT compression the bond dimension is rather small, because only few exponentials contribute.
-However, a peak in the bond dimension can be seen in the vicinity of the ground state crossover, indicated by the dashed lines.
-Here, the singlet ground state at half filling crosses with a doublet that has one (three) electrons at small (large) values of $\mu/U$.
-The picture becomes clearer when analyzing the bond dimension of the local two-particle Green’s function $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}(\tau_{1},\tau_{2},\tau_{3})=-\langle\mathcal{T}\hat{c}_{1,\uparrow}(\tau_{1})\hat{c}_{1,\uparrow}^{\dagger}(\tau_{2})\hat{c}_{1,\uparrow}(\tau_{3})\hat{c}_{1,\uparrow}^{\dagger}\rangle$ in Fig.
-1 (b) with $\epsilon=10^{-14}$ at $\beta=50$ and $R=6$.
-Here, for three imaginary times more exponentials contribute giving rise to a smoother picture with larger bond dimensions.
-Thus, a sharp peak can be seen as a remnant of the QPT.
-Increasing temperature ($\beta=30$) in Fig.
-1(d) leads to the expected smearing in the vicinity of the transition.
-In Fig.
-1(c), the value of the spin susceptibility $\chi_{S}$ (see Appendix for definition) is shown, clearly distinguishing the two “phases”.
-In the Supplemental Material [40], we show that this analysis also applies to other correlators (including Matsubara frequency dependence) and temperatures, and also to crossings of the first excited states.
+1(a), the value of the spin susceptibility $\chi_{S}$ (definition in App.
+E) is shown, clearly distinguishing the two “phases”.
+Fig.
+1(c) plots the bond dimension as time scale entanglement measure over the imaginary time scale.
+In App.
+A and the Supplemental Material [49], we show that this analysis also applies to other correlators (including Matsubara frequency dependence) and temperatures, and also to crossings of the first excited states.
 We additionally demonstrate how to proceed with low-accuracy data by applying higher QTT cutoffs.
 VI Four-site Hubbard ring
-Next, let us consider the periodic 4-site Hubbard ring with nearest neighbor hopping only, i.e., $N=4,t^{\prime}=0$ in Eq.
-(IV).
+We consider the periodic 4-site Hubbard ring with nearest neighbor hopping only, i.e., $N=4,t^{\prime}=0$, $t_{\pm(N-1)}=t$, in Eq.
+(5).
 In Fig.
-2(a), the bond dimension of the QTT representation of the two-particle Green’s function $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}(\tau_{1},\tau_{2},\tau_{3})$ is shown in the $\mu/U-U$-plane.
-Crossings of the ground state energy are indicated by red dots.
-For not too low values of $U$ it can be seen that the dots lie precisely on top of the peaks of the bond dimension.
-This becomes also visible in Fig.
-2(e), where the slice for $U=4$ is shown and the remnants of the QPT (red dotted lines) can be seen as sharp peaks in $D_{\mathrm{max}}$.
-For lower values of $U$ the bond dimension peaks deviate from the crossings in the eigenspectrum of the Hamiltonian (red dashed lines) and become broader.
+2(a), $D_{\mathrm{max}}$ of the QTT representation of the two-particle Green’s function $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}(\tau_{1},\tau_{2},\tau_{3})$ is shown in the $\mu/U-U$-plane.
+Crossings of the ground states are indicated by red dots.
+For not too low values of $U$ the dots lie precisely on top of the peaks of $D_{\mathrm{max}}$.
+In Fig.
+2(e), the slice for $U=4$ is shown.
+For lower values of $U$ the $D_{\mathrm{max}}$ peaks deviate from the crossings in the eigenspectrum of the Hamiltonian (red dashed lines) and become broader.
 This is because for small $U$ and $\beta=50$ excited states become thermally accessible.
-Consequently, when decreasing $T$, the region with large bond dimensions shifts to smaller $U$’s, see Fig.
+Consequently, when decreasing $T$, the region with large $D_{\mathrm{max}}$ shifts to smaller $U$’s, see Fig.
 2(f), and for $T\rightarrow 0$ again alienates with the eigenvalue crossings.
 The spin susceptibility $\chi_{S}$ in Fig.
-2(c) and the fidelity susceptibility $\chi_{F}$ (in Fig.
-2(d) (see Appendix C for its definition) show a qualitatively similar behavior as the bond dimension: at large values of $U$ the ground state crossings are marked by a maximum, while for small values of $U$ both also show larger values, absent of an eigenvalue crossing.
+2(c) and, e.g., the fidelity susceptibility $\chi_{F}$ (in Fig.
+2(d) (definition in App.
+C) show a qualitatively similar behavior as $D_{\mathrm{max}}$: at large values of $U$ the ground state crossings are marked by a maximum, while for small values of $U$ both show larger values, without an eigenvalue crossing.
 Further, when investigating entanglement in Fig.
-2(b), also a similar structure is seen, where the white dashed line indicates the quantum Fisher information $F_{Q}=1$ (see Appendix for definition).
-Hence, the peaks in the bond dimension for low $U$ in Fig.
+2(b), also a similar structure is seen, where the white dashed line indicates the quantum Fisher information $F_{Q}=1$ (definition in App.
+F).
+Hence, the peaks in $D_{\mathrm{max}}$ for low $U$ in Fig.
 2(a) seem to be associated with some thermal crossover between different regimes.
-Further evidence of this thermal crossover of regimes seen by QTTD can be seen in Fig.
+Further evidence of this thermal crossover seen by QTTD is seen in Fig.
 2(f), where the QTT bond dimension of $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}$ is shown for various temperatures.
-The dashed lines indicate where $F_{Q}=1$ and are located at the maxima of the different curves.
-Interestingly, also the derivative of the quantum Fisher information $\frac{dF_{Q}}{dU}$ (and other entanglement measures - see SM [40]) as well as that of the double occupancy show a maximum in close vicinity of this point.
-Additionally, also the maximum and inflection point of the spin susceptibility lie in proximity to this point and show a similar temperature behavior.
-We refer the reader to the SM [40] for further details and figures on all of the properties discussed.
-Figure 2: Four-site Hubbard ring with nearest-neighbor hopping only: (a) $D_{\mathrm{max}}$ of the QTT for $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}(\tau_{1},\tau_{2},\tau_{3})$ at $\beta=50$, $\epsilon=10^{-14}$ and $R=6$ in comparison to (b) the quantum Fisher information $F_{Q}$, where the white dashed line indicates $F_{Q}=1$ , (c) spin susceptibility $\chi_{S}$ and (d) fidelity susceptibility.
+The dashed lines indicate $F_{Q}=1$ and are located at the maxima of the different curves.
+Interestingly, also the derivative $\frac{dF_{Q}}{dU}$ (and other entanglement measures - see SM [49]) as well as that of the double occupancy show a maximum in close vicinity of this point.
+Additionally, also the maximum and inflection point of the spin susceptibility lie in proximity to this point and show a similar temperature behavior (further details in SM [49]).
+Figure 2: Four-site Hubbard ring with nearest-neighbor hopping only: (a) $D_{\mathrm{max}}$ of the QTT for $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}(\tau_{1},\tau_{2},\tau_{3})$ at $1/T=\beta=50$, $\epsilon=10^{-14}$ and $R=6$ in comparison to (b) the quantum Fisher information $F_{Q}$, where the white dashed line indicates $F_{Q}=1$ , (c) spin susceptibility $\chi_{S}$ and (d) fidelity susceptibility.
 (e) and (f): $U=4$ and $\mu=U/2$ slice of (a), where red dashed lines in (e) indicate QPTs and dashed lines in (f) indicate $F_{Q}=1$.
@@
 VII Four-site Hubbard ring with next-nearest neighbor hopping
-Next, as an indicator of time scale entanglement in the system, the bond dimension of the QTT representation of the two-particle correlator $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}$ of the periodic 4-site Hubbard ring with next-nearest neighbor hopping $t^{\prime}$ at $U=10$, $\beta=50$ is studied in the $t^{\prime}-\mu/U$-plane [41].
-In contrast to the previous examples, we apply larger QTT cutoffs $\epsilon$.
+We consider $D_{\mathrm{max}}$ of $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}$ of the periodic 4-site Hubbard ring with next-nearest neighbor hopping $t^{\prime}$ at $U=10$, $\beta=50$ in the $t^{\prime}-\mu/U$-plane [50].
+Contrary to the previous examples, we apply larger QTT cutoffs $\epsilon$.
 In Fig.
@@
 Still, the different ground state crossings (indicated by red dots) between s(inglet), d(oublet) and q(uartet) states lie precisely on top of the peaks of $D_{\mathrm{max}}$.
-This becomes even better visible in Figs.
-3(b)-(c), where different $\mu/U$-slices are shown for a different QTT cutoff of $\epsilon=10^{-8}$.
-Here, the sharp peaks of the bond dimension are again precisely located at the QPTs demonstrating how well the proposed QTTD method works by producing stable results for different QTT cutoffs ($\epsilon=10^{-5}$ in (a) vs.
+In Figs.
+3(b)-(c), different $\mu/U$-slices are shown for a lower QTT cutoff of $\epsilon=10^{-8}$, demonstrating how well the proposed QTTD works by producing stable $D_{\mathrm{max}}$ peaks for different QTT cutoffs ($\epsilon=10^{-5}$ in (a) vs.
 $\epsilon=10^{-8}$ in (b)-(c)).
-This shows that the time scale entanglement diagnosed by the growth of the bond dimension $D_{\mathrm{max}}$ is system-inherent and visible across all studied correlators (see SM [40] additional plots).
-In App.
-B, a short recipe on how QTT cutoffs can be efficiently used, to deal with low-accuracy data is presented and in the SM [40], we comment on crossings that are correctly diagnosed by QTTD for which $\chi_{F}$ fails to achieve the same.
+This shows that time scale entanglement diagnosed by the growth of $D_{\mathrm{max}}$ is system-inherent and visible across all studied correlators (see SM [49] additional plots).
+In the SM [49], we comment on crossings that are correctly diagnosed by QTTD for which $\chi_{F}$ fails to achieve the same.
 Figure 3: Four-site Hubbard ring with next nearest-neighbor hopping: Red dots (dotted lines) indicate crossings of s(inglet), d(oublet) and q(uartet) ground states.
@@
 VIII Single-impurity Anderson model
-We examine the single-impurity Anderson model (SIAM), which features a magnetic impurity in a metallic bath and shows, among others, a crossover between local moment and Kondo regime.
+The single-impurity Anderson model (SIAM) features a magnetic impurity in a metallic bath and shows a crossover between local moment and Kondo regime.
 For numerical calculations, the continuous bath is discretized leading to the following Hamiltonian
@@
 $+\sum_{\ell=1}^{3}E_{\ell}\hat{c}_{\ell}^{\dagger}\hat{c}_{\ell}+\sum_{\ell=1}^{3}(V_{\ell}\hat{d}_{\sigma}^{\dagger}\hat{c}_{\ell}+\mathrm{h.c.}).$ (6)
-Here, $\varepsilon_{0}$ represents the impurity site energy, $\mu$ is the chemical potential, and $U$ is the on-site Coulomb interaction on the impurity site.
-The annihilation (creation) operator on the impurity site is denoted by $\hat{d}_{\sigma}^{({\dagger})}$, and the number operator is $\hat{n}_{\sigma}=\hat{d}_{\sigma}^{\dagger}\hat{d}_{\sigma}$.
-The density of states of the bath is assumed to be semicircular, and the parameters $V_{\ell}$ and $E_{\ell}$ denote the hybridization strength and energy levels of the discretized bath, respectively.
-The continuous bath is discretized into three levels, as detailed in the SM [40].
-Figure 4(a) illustrates the bond dimension of the QTT representation of the two-particle Green’s function $G^{\uparrow\uparrow\uparrow\uparrow}_{\mathrm{imp}}(\tau_{1},\tau_{2},\tau_{3})$ at the impurity site for $\beta=100$, plotted against the model parameters $U$ and $V$, where $V$ denotes the hybridization strength in the original continuous model.
+$\varepsilon_{0}$ represents the impurity site energy, $\mu$ is the chemical potential, and $U$ the on-site interaction.
+The annihilation (creation) operator on the impurity site is $\hat{d}_{\sigma}^{({\dagger})}$, and $\hat{n}_{\sigma}=\hat{d}_{\sigma}^{\dagger}\hat{d}_{\sigma}$.
+The bath density of states is assumed to be semicircular, and the parameters $V_{\ell}$ and $E_{\ell}$ denote the hybridization strength and energy levels of the discretized bath, respectively.
+The bath is discretized into three levels, as detailed in the SM [49].
+Figure 4(a) illustrates $D_{\mathrm{max}}$ of the QTT representation of $G^{\uparrow\uparrow\uparrow\uparrow}_{\mathrm{imp}}(\tau_{1},\tau_{2},\tau_{3})$ at the impurity site for $\beta=100$, plotted against the model parameters $U$ and $V$, where $V$ denotes the hybridization strength in the original continuous model.
 In Fig.
 4(b), the spin susceptibility $\chi_{S}$ is depicted.
-The red dots mark the Kondo temperature $T_{K}=V\sqrt{U}e^{-\pi U/8V^{2}}$ estimated via poor man’s scaling [42, 43], which delineates the boundary between the Kondo and local moment regimes, as also reflected in the spin susceptibility behavior.
-The bond dimension has a broad peak located in the vicinity of the Kondo temperature, diagnosing the thermally driven Kondo to local moments regime crossover.
-Figure 4: (a) Bond dimension $D_{\mathrm{max}}$ of the two-particle Green’s function of the single-impurity Anderson model at $\beta=100$ in the $U$-$V$ plane, where the red dots indicate $T_{K}(U,V)=1/\beta$.
+The red dots mark the Kondo temperature $T_{K}=V\sqrt{U}e^{-\pi U/8V^{2}}$ estimated via poor man’s scaling [51, 52], which delineates the boundary between Kondo and local moment regimes, as also reflected in the spin susceptibility behavior.
+$D_{\mathrm{max}}$ has a broad peak located in the vicinity of the Kondo temperature, diagnosing the thermally driven Kondo to local moments regime crossover.
+Figure 4: (a) $D_{\mathrm{max}}$ of the two-particle Green’s function of the SIAM at $\beta=100$ in the $U$-$V$ plane [red dots indicate $T_{K}(U,V)=1/\beta$].
 (b) Spin susceptibility $\chi_{S}$ indicating the different regimes.
@@
 Since their recent invention, QTTs have primarily been applied as a compression tool in numerical calculations to overcome severe memory bottlenecks.
-In this work, we have presented an entirely novel approach on diagnosing phase transitions and crossovers with QTTs.
-We have demonstrated that this new quantics tensor train diagnostics approach is able to identify QPTs and thermal regime crossovers and distinguish between these two by studying various one-dimensional Hubbard rings and the SIAM.
-Moreover, by showing results for different correlators we have identified time scale entanglement to be system-inherent showing maxima at phase transitions and crossovers and, additionally, have pointed out that this method is also able to deal with low-accuracy data by varying the QTT cutoff.
-For the four-site ring with next-nearest neighbor hopping, QTTD correctly diagnosed ground state crossings that the fidelity susceptibility failed to capture.
-The strong point of the QTTD is that it can be used universally for zero and finite temperature and multiple correlation functions – it is not limited to ground state methods and does not require specific measures or susceptibilities.
+In this work, we present a novel QTTD approach, diagnosing time-scale entanglement from the QTT bond dimension.
+We further demonstrate that QTTD is actually able to identify thermal crossovers, ground state changes, and presumably QPTs, and to distinguish between these phenomena.
+Studying various one-dimensional Hubbard rings and the SIAM, we show that time-scale entanglement for all correlators becomes maximal at QPTs and crossovers.
+This is even true for the one-particle Green’s function whose static, time-independent part does not signal a second order phase transition when approaching it from the symmetry-unbroken side.
+In the Green’s function dynamics, the phase transition is however encoded in its time-scale entanglement.
+Additionally, we point out that QTTD is able to deal with low-accuracy data by varying the QTT cutoff.
+Comparing to conventional entanglement measures, we show that QTTD correctly diagnoses ground state crossings that the fidelity susceptibility fails to capture.
+The strong point of QTTD is that it can be used universally for zero and finite temperature and multiple correlation functions—it is not limited to ground state methods and does not require specific measures or susceptibilities.
 Further, QTTD offers a new perspective and deeper understanding on the inherent time and length scale entanglement at phase transitions.
+It can provide such a fresh perspective also for other physical phenomena.
 Acknowledgements.
@@
 For open access purposes, the authors have applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission.
-Calculations have been partly performed on the Vienna Scientific Cluster (VSC).
+Calculations have been partly performed using Austrian Scientific Computing (ASC).
 H.S.
@@
 JPMJFR2232, Japan.
-H.I.
-and H.S.
-were supported by JST PRESTO Grant No.
-JPMJPR2012, Japan.
 Appendix A Appendix A: QTTD and multi-point correlators
 Let us consider an arbitrary n-point correlator function $\mathcal{G}(\bm{\tau})=-\langle\mathcal{T}\hat{O}_{1}(\tau_{1})\hat{O}_{2}(\tau_{2})...\hat{O}_{n}(0)\rangle$ with $\bm{\tau}=(\tau_{1},\tau_{2},...,\tau_{n-1}),\tau_{i}\in[0,\beta)$ and fermionic or bosonic operators $\hat{O}_{i}$.
-We express the correlator in terms of its Lehmann representation
-$\mathcal{G}(\bm{\tau})$ $=-\frac{1}{Z}\sum_{m_{1}}e^{-\beta\epsilon_{m_{1}}}\sum_{m_{2}...m_{n}}e^{-\tau_{1}(\epsilon_{m_{2}}-\epsilon_{m_{1}})}e^{-\tau_{2}(\epsilon_{m_{3}}-\epsilon_{m_{2}})}...e^{-\tau_{n-1}(\epsilon_{m_{n}}-\epsilon_{m_{n-1}})}$
-×$\times\langle m_{1}|\hat{O}_{1}|m_{2}\rangle\langle m_{2}|\hat{O}_{2}|m_{3}\rangle...\langle,m_{n}|\hat{O}_{n}|m_{1}\rangle$ (7)
+For $\tau_{1}>\tau_{2}>\ldots\tau_{n-1}$, we express the correlator in terms of its Lehmann representation:
+$\mathcal{G}(\bm{\tau})$ $=-\frac{1}{Z}\sum_{m_{1}\ldots m_{n}}e^{-\tau_{1}(\epsilon_{m_{2}}-\epsilon_{m_{1}})}e^{-\tau_{2}(\epsilon_{m_{3}}-\epsilon_{m_{2}})}\cdots e^{-\tau_{n-1}(\epsilon_{m_{n}}-\epsilon_{m_{n-1}})}$
+$\times e^{-\beta\epsilon_{m_{1}}}\langle m_{1}|\hat{O}_{1}|m_{2}\rangle\langle m_{2}|\hat{O}_{2}|m_{3}\rangle\cdots\langle m_{n}|\hat{O}_{n}|m_{1}\rangle$ (7)
 with the eigenstates $|m_{i}\rangle$ of the Hamiltonian and we set the ground state energy $\epsilon_{\mathrm{GS}}=0$ in the following.
@@
 Since the QTT representation of an exponential has a bond dimension of one and the bond dimension of the sum of two QTTs is bound by the sum of the bond dimensions of the individual QTTs, we expect three-point correlators to have larger QTT bond dimensions than two-point correlators due to additional imaginary time dependence and more dominating contributions in the sum.
-Therefore, higher-point correlators are also expected to show more pronounced peaks at phase transitions and crossovers, which is confirmed by the results in the paper.
+Therefore, higher-point correlators are also expected to show more pronounced peaks at phase transitions and crossovers (cf.
+Fig.
+1 and Fig.
+5).
 The reasoning that phase transitions and crossovers are connected to maxima in the bond dimensions of correlators follows the same arguments as in the main text.
 Let us emphasize that this derivation is valid for arbitrary (non-zero) correlators.
-Since Fourier transformation can be represented by a low-rank matrix product operator applied to the correlator QTT, imaginary frequency counterparts to the imaginary times correlator are expected to display the same maxima in the bond dimension, which is supported by the results in the Supplemental Material.
+Since Fourier transformation can be represented by a low-rank matrix product operator [38, 53] applied to the correlator QTT, imaginary frequency counterparts to the imaginary times correlator are expected to display the same maxima in the bond dimension, which is supported by the results in the Supplemental Material.
 Therefore, arbitrary correlators that are available from computations can be used in QTTD to diagnose phase transitions and crossovers in systems.
@@
 Further evidence and support of this property is provided in the SM [49].
+Figure 5: QTT bond dimension $D_{\mathrm{max}}$ of different imaginary-time correlation functions for the Hubbard dimer as a function of electronic repulsion $U$ and chemical potential $\mu/U$: (a) $D_{\mathrm{max}}$ for the one-particle Green’s function $G_{11}^{\uparrow\uparrow}$ at temperature $T=\beta^{-1}=1/50$, and (b) for the two-particle Green’s function $G_{1111}^{\uparrow\uparrow\uparrow\uparrow}$ at $T=1/30$.
+Red dashed lines mark crossings of the ground state of the model.
+These crossings can be diagnosed as a sharp maximum in $D_{\mathrm{max}}$.
 Appendix C Appendix C: Fidelity susceptibility
-Following Ref. 44, we define the fidelity susceptibility [45, 46, 47, 48, 49] at finite temperatures as
-[…]
+Following Ref. 54, we define the fidelity susceptibility [55, 56, 57, 58, 59] at finite temperatures as
+$\chi_{F}=\int_{0}^{\frac{\beta}{2}}d\tau\left[\langle\hat{H}_{1}(\tau)\hat{H}_{1}\rangle-\langle\hat{H}_{1}\rangle^{2}\right]\tau$ (9)
 with
@@
 The local spin susceptibility $\chi_{S}$ is defined in the following way [52]
-[…]
+$\chi_{s}=\int_{0}^{\beta}d\tau\langle\hat{S}_{z}(\tau)\hat{S}_{z}(0)\rangle,$ (11)
 where $\hat{S}_{z}=(\hat{n}_{\uparrow}-\hat{n}_{\downarrow})/2$ describes the local magnetization on a single site. Similarly, the local charge susceptibility is defined as follows
-[…]
+$\chi_{c}=\int_{0}^{\beta}d\tau\langle\hat{n}(\tau)\hat{n}(0)\rangle,$ (12)
 with $\hat{n}=(\hat{n}_{\uparrow}+\hat{n}_{\downarrow})/2$.